OpenStudy (anonymous):

Consider the prism shown below. a.Draw a net for the prism and label all dimensions. b.Use the net to find the surface area of the prism

OpenStudy (anonymous):

@satellite73 @e.mccormick @thomaster Help? Yes, no, maybe? Yes. :)

OpenStudy (e.mccormick):

Well, you know what the net is?

OpenStudy (e.mccormick):

http://math8geometry.wikispaces.com/How+To+Draw+A+Net+For+A+Cylinder+%26+Triangle Once you have made the net, the surface area is the total area of all the basic shapes.

OpenStudy (anonymous):

I got 456m. But someone said it was 626 squared meters. I'm quite confused.

OpenStudy (anonymous):

@e.mccormick

OpenStudy (e.mccormick):

 I get 456. I don't know where they got the extra from.

OpenStudy (e.mccormick):

(10+8+6)17+6*8=24*17+48=408+48=456